OpenStudy (anonymous):

which is the polynomial written as the product of two binomials? 5x^3-10x^2+x-2

OpenStudy (anonymous):

5x^3-10x^2+x-2 = x ( 5x^2 + 1) - 2 ( 5x^2 +1) = ( 5x^2 + 1) ( x -2)